1/2 multiply by 2 1/10
1
Expert's answer
2020-03-11T14:15:37-0400

"1/2 × 2 1/10 = \frac{1} {2} × \frac{21} {10} = \frac{21} {20} = 1.05 = 1 \frac{1} {20}"